simplify: sqrt (90)
a) 30
b) 3 sqrt (10)
c) 10 srt (3)
d) 9 sqrt (3)

Found 2 solutions by CharlesG2, texttutoring:
Answer by CharlesG2(834)

simplify: sqrt (90)
a) 30
b) 3 sqrt (10)
c) 10 srt (3)
d) 9 sqrt (3)
sqrt or square root means you figure out what number times itself equals the number you are finding the square root of
for example the square root of 9 is 3 because 3^2 = 3*3 = 9
the square root of 10 can not be easily done so we leave that sqrt(10)
sqrt(90) = sqrt(9 * 10) = 3sqrt(10)
the answer is b

Answer by texttutoring(324)

sqrt(90) = sqrt(9*10) = sqrt(9) * sqrt(10) = 3 * sqrt(10)

The answer is b.